OpenStudy (anonymous):

interval notation 5x+10>50 I got x=8 how do I write it in interval notation (-inf,8)

OpenStudy (shubhamsrg):

you got x>8 and not x=8,,this little thing helps ?

OpenStudy (anonymous):

yes I got it but how do I write is as a interval notation (-\infty,8)\]

OpenStudy (anonymous):

\[(-\infty,8)\]

OpenStudy (shubhamsrg):

does 9 fall in (-inf,8) ?

OpenStudy (anonymous):

hun, I am lost what does 9 have to do with this

OpenStudy (shubhamsrg):

i just randomly picked a no..according to x>8,,9 should be a solution,,but does your interval include 9 ?

OpenStudy (anonymous):

no i don't think so

OpenStudy (anonymous):

do i write it as \[(-\infty,8) or (8,-\infty) or ( 8,\infty)\]

OpenStudy (shubhamsrg):

which one according to you contains 9,,or 10..etc ?

OpenStudy (anonymous):

I have no idea where this 9 even came into this help

OpenStudy (kropot72):

[infinity,8) The above interval notation means the interval of all the real numbers between infinity and 8, including infinity but excluding 8.

OpenStudy (shubhamsrg):

lol..i just randomly picked a no. hoping to make things simpler..hmm,,nevermind.. ans will be (8,inf)

OpenStudy (anonymous):

why is it \[(8,\infty)\] and not the other way around

OpenStudy (anonymous):

I am so lost with interval notation I can not make heads or tails of it

OpenStudy (anonymous):

thanks I will look at it right now

OpenStudy (kropot72):

http://www.mathwords.com/i/interval_notation.htm The above link gives detailed information on the format for interval notation.

OpenStudy (anonymous):

-5x-12<8 I got -4 so i put it in interval notation (-4,inf)

OpenStudy (shubhamsrg):

nops..again wrong..should have been (-inf,-4)

OpenStudy (anonymous):

YYYYYYYYYYY

OpenStudy (anonymous):

wow, I sound like a big baby

OpenStudy (shubhamsrg):

see when it was x>8, it meant x can take values 8.8 ,9 ,9.7 ,100 ,12678, etc,,till infinty so we write it as (8,inf) signifying from 8 to infinity ! now can you see why that one is wrong..

OpenStudy (anonymous):

oh! I am so slow thanks

OpenStudy (anonymous):

so because the x is larger the number is first and positive

OpenStudy (shubhamsrg):

hmm..

OpenStudy (anonymous):

guess I am still lost

OpenStudy (shubhamsrg):

try this one 2x + 6 > x -7 what should be the ans ?